the mountain springs water company has two departments purifying and bottling, The bottling department received 61090 liters from the purifying department. during the period the bottling department completed 58410 liters, including 2590 liters of work in process at the beginning of the period. The ending work in process was 5270 liters how many liters were started and completed during the period
Answer #1

Correct answer----------Liters started and completed = 55,820

Working

Reconciliation of Units
A Opening WIP                2,590
B Introduced              61,090
C=A+B TOTAL              63,680
D Transferred              58,410
E=C-D Closing WIP                5,270

Started and completed = Total transferred-units in beginning

Started and completed= 58410-2590= 55820 Units